### v1

@@ Abstract
Nonsmooth nonconvex-concave minimax problems have attracted significant attention due to their wide applications in many fields. In this paper, we consider a class of nonsmooth nonconvex-concave minimax problems on Riemannian manifolds. Owing to the nonsmoothness of the objective function, existing minimax manifold optimization methods cannot be directly applied to solve this problem. We propose a manifold proximal gradient descent ascent (MPGDA) algorithm for solving the problem. At each iteration, the proposed algorithm alternately performs one or multiple manifold proximal gradient descent steps and a proximal ascent step. We prove that the MPGDA algorithm can find an $\varepsilon$-game-stationary point and an $\varepsilon$-optimization-stationary point of the considered problem within $\mathcal{O}(\varepsilon^{-3})$ iterations. Numerical experiments on fair sparse PCA and sparse spectral clustering are conducted to demonstrate the advantages of the MPGDA algorithm.
@@ 1 Introduction
where $\mathcal{M}$ is a Riemannian manifold embedded in a finite-dimensional Euclidean space $\mathbb{E}_{1}$, and $S$ is a convex set in another finite-dimensioanl Euclidean space $\mathbb{E}_{2}$, $f:\mathbb{E}_{1}\times\mathbb{E}_{2}\to\mathbb{R}$ is a continuously differentiable function and is concave with respect to y, $h:\mathbb{E}_{1}\to\mathbb{R}$ and $g:\mathbb{E}_{2}\to\mathbb{R}$ are some proper closed convex (possibly nonsmooth) functions. Throughout this paper, we adopt the following assumptions for problem (1) .
@@
$(i)$ The manifold $\mathcal{M}$ is compact.
@@
Problem (1) has many important applications in machine learning and signal processing. Below, we present two representative examples and refer the interested readers to chenProximalGradientMethod2020 ; BriefIntroductiontoManifoldOptimization for more examples.
@@
Fair Sparse Principle Component Analysis (FSPCA). The classical PCA Hotelling1933Analysis is one of the most widely used dimensional reduction techniques, which minimizes the total reconstruction error over the whole samples. However, in real-life applications, the data samples come from diverse classes and PCA may exhibit a higher reconstruction error for certain classes than the others. To address this disparity, the fair PCA is proposed samadiPriceFairPCA . Suppose the $m$ data samples belong to $n$ classes, and each class $i$ corresponds to $A_{i}\in\mathbb{R}^{m_{i}\times d}$ with $d$ being the dimension of the samples, $m_{i}$ being the number of samples in the $i$-th class, and $\sum_{i=1}^{n}m_{i}=m$. Then the fair PCA can be formulated as:
@@
where $r<d$ is the number of principle components, $\mathrm{Tr}(\cdot)$ represents the trace and $\mathrm{St}(d,r):=\{X\in\mathbb{R}^{d\times r}|X^{\top}X=I_{r}\}$ with $I_{r}$ denoting the $r\times r$ identity matrix. Recently, the fair sparse PCA is further proposed to promote sparsity in the principle components babuFairPrincipalComponent2023 as follows:
@@
where $\mu>0$ is a weighting parameter and $\left\|X\right\|_{1}:=\sum_{i,j}\left|x_{ij}\right|$ is the $\ell_{1}$-norm of the matrix $X$. In view of (Beck2017First, , Example 4.10), problem (3) can be equivalently reformulated into
@@
Sparse Spectral Clustering (SSC). The SSC problem aims to partition $N$ data samples into $p$ groups such that similar data points are clustered together. In spectral clustering, a symmetric affinity matrix $W=[W_{ij}]_{N\times N}$ is constructed, where $W_{ij}\geq 0$ quantifies the pairwise similarity between samples $a_{i}$ and $a_{j}$. To promote sparsity and interpretability, the sparse spectral clustering framework is introduced in 2016ConvexSparseSpectralClustering , as follows
@@
In recent years, nonconvex-concave (NC-C) minimax problems have received a tremendous amount of attention as they have found broad applications in diverse areas. Many studies focus on NC-C problems in the setting where both the constraints on $x$ and $y$ are closed convex sets and numerous methods have been developed, see, for example, linGradientDescentAscent ; linOptimalAlgorithmsMinimax ; nouiehedSolvingClassNonConvex ; ostrovskiiEfficientSearchFirstOrder2021 ; panEfficientAlgorithmNonconvexlinear2021a ; 2022Zeroth ; thekumparampilEfficientAlgorithmsSmooth ; xuUnifiedSingleloopAlternating2023 ; yangFasterSingleloopAlgorithms ; zhangSingleLoopSmoothedGradient . In contrast, there are relatively few works studying NC-C minimax problems in the setting where the constraint on $x$ is a Riemannian submanifold and the constraint on $y$ is a closed convex set. Among these works, Huang and Gao huangGradientDescentAscent2023 propose a Riemannian gradient descent-ascent algorithm for a smooth nonconvex-strongly concave (NC-SC) minimax problems and establish its iteration complexity of $\mathcal{O}(\varepsilon^{-2})$ for achieving $\varepsilon$-stationarity. Later, Xu et al. xuEfficientAlternatingRiemannian2024 present an alternating Riemannian/projected gradient descent-ascent (ARPGDA) algorithm with an iteration complexity of $\mathcal{O}(\varepsilon^{-3})$ for finding an $\varepsilon$-stationary point of smooth Riemannian nonconvex-linear (NC-L) minimax problem. Very recently, for problem (1) with additional assumptions that $h$ is smooth (but not necessarily convex) and $f(x,y)$ is linear with respect to $y$ for fixed $x$, the work xuRiemannianAlternatingDescent2024 extend the technique in and develop algorithms that can reach $\varepsilon$-stationary points within $\mathcal{O}(\varepsilon^{-3})$ iterations. It is worth noting that none of the methods in the aforementioned literature huangGradientDescentAscent2023 ; xuRiemannianAlternatingDescent2024 ; xuEfficientAlternatingRiemannian2024 can directly tackle problem (1) due to the following two reasons: (i) the convex function $h$ may be nonsmooth, and (ii) the function $f(x,y)$ is neither linear nor strongly concave with respect to $y$ for fixed $x$. Based on the needs of applications and the above background, it is highly desirable to develop efficient methods with theoretical convergence guarantees for solving the Riemannian nonsmooth NC-C minimax problem (1).
@@
The main contributions of this paper are summarized below. First, we introduce the notions of optimization stationarity point as well as game stationarity for problem (1) and carefully study their relationships. Second, we propose a manifold proximal gradient descent ascent method for solving problem (1). At each iteration, this method performs one or multiple manifold proximal gradient descent steps to update $x$ and one proximal ascent step to update $y$. Third, we prove that the proposed method can find an $\varepsilon$-game-stationary point and an $\varepsilon$-optimization-stationary point within $O(\varepsilon^{-3})$ iterations. This iteration complexity result matches the state-of-the-art complexity analysis for single-loop algorithms in solving general nonsmooth NC-C optimization problems, where the constraint on $x$ is a compact convex set rather than a Riemannian manifold. Finally, we present numerical results on FSPCA and SSC to demonstrate the effectiveness of our proposed method.
@@
The rest of the paper is organized as follows. Section 2 presents preliminaries that will be used in this paper. Section 3 discusses the optimality condition of problem (1). Section 4 introduces the proposed MPGDA method and establish its iteration complexity. Section 5 reports the numerical experiments.
@@ 2 Preliminaries
[…]
@@ 2.1 Riemannian submanifold
(boumal2023intromanifolds, , Definition 3.10) Let $\mathbb{E}$ be a Euclidean space of dimension $d$. A nonempty subset $\mathcal{M}$ of $\mathbb{E}$ is called an embedded submanifold of $\mathbb{E}$ of dimension $n$ if either $n=d$ and $\mathcal{M}$ is open in $\mathbb{E}$ or $n=d-k$ for some $k\geq 1$ and, for each $x\in\mathcal{M}$, there exists a neighborhood $U$ of $x$ in $\mathbb{E}$ and a smooth function $\psi:U\to\mathbb{R}^{k}$ such that:
@@
Let $\mathcal{M}$ be an embedded submanifold of $\mathbb{E}$. The tangent space $T_{x}\mathcal{M}$ to $\mathcal{M}$ at $x\in\mathcal{M}$ is identified with $\mathbb{E}$ when $\mathcal{M}$ is open in $\mathbb{E}$; otherwise, $T_{x}\mathcal{M}=\mathrm{ker}(\mathrm{D}\psi(x))$ , where $\psi$ is a local defining function for $\mathcal{M}$ at $x$ (boumal2023intromanifolds, , Theorem 3.15). The embedded submanifold $\mathcal{M}$ is called a Riemannian submanifold of $\mathbb{E}$, when equipped with the Riemannian metric induced by the ambient space $\mathbb{E}$. Namely, for each $x\in\mathcal{M}$, the inner product on $T_{x}\mathcal{M}$ is defined as $\langle u,v\rangle_{x}:=\langle u,v\rangle$ for $u,~{}v\in T_{x}\mathcal{M}$. The normal space $N_{x}\mathcal{M}$ of a Riemannian submanifold $\mathcal{M}$ of $\mathbb{E}$ is the orthogonal complement of the tangent space $T_{x}\mathcal{M}$ in $\mathbb{E}$ (boumal2023intromanifolds, , Definition 5.48).
The tangent bundle of $\mathcal{M}$ is the disjoint union of its tangent spaces : $T\mathcal{M}:=\{(x,v)\mid x\in\mathcal{M},~{}v\in T_{x}\mathcal{M}\}$. A retraction on a manifold $\mathcal{M}$ is a smooth mapping $R$ from the tangent bundle $T\mathcal{M}$ onto $\mathcal{M}$ such that, for any $(x,v)\in T_{x}\mathcal{M}$, its restriction $R_{x}:T_{x}\mathcal{M}\to\mathcal{M}$ satisfies (i) $R_{x}(0)=x$ and (ii) $\mathrm{D}R_{x}(0)[v]=v$ (boumal2023intromanifolds, , Definition 3.47). The concept of a retraction offers a practical link between manifold and its tangent bundle. The following properties of retractions are needed for convergence analysis in this paper.
@@
[…]
@@ 2.2 Generalized subdifferential
For a fucntion $\varphi:\mathbb{E}\to\mathbb{R}$, the usual one-sided directional derivative of $\varphi$ at $x\in\mathbb{E}$ along the direction $v\in\mathbb{E}$ is defined by
@@
In contrast to the usual one-sided directional derivative, the (Clarke) generalized directional derivative dose not presuppose the existence of any limit (1983Optimization, , Section 2.1). Specifically, the (Clarke) generalized directional derivative of $\varphi$ at $x\in\mathbb{E}$ along the direction $v\in\mathbb{E}$ is defined as
@@
The function $\varphi$ is said to be regular at $x$, provided that for all $v\in\mathbb{E}$, the usual one-sided directional derivative $\varphi^{\prime}(x;v)$ exists and $\varphi^{\prime}(x;v)=\varphi^{\circ}(x;v)$. If $\varphi_{1}$ and $\varphi_{2}$ are regular at $x$, then $\partial(\varphi_{1}+\varphi_{2})(x)=\partial\varphi_{1}(x)+\partial\varphi_{2}(x)$ (1983Optimization, , Definition 2.3.4). If $\varphi$ is a continuously differentiable function or a real-valued convex function, then it is regualr at any $x\in\mathbb{E}$ (1983Optimization, , Proposition 2.3.6). In addition, the indication function $\partial\delta_{\mathcal{M}}(x)$ of a Remannian submanifold $\mathcal{M}$ in $\mathbb{E}$ is regular at any $x\in\mathcal{M}$ (1983Optimization, , Theorem 2.4.7) and $\partial\delta_{\mathcal{M}}(x)=N_{x}\mathcal{M}$ (1983Optimization, , Proposition 2.4.12).

### v2

@@ Abstract
Nonsmooth nonconvex-concave minimax problems have attracted significant attention due to their wide applications in many fields. In this paper, we consider a class of nonsmooth nonconvex-concave minimax problems on Riemannian manifolds. Owing to the nonsmoothness of the objective function, existing minimax manifold optimization methods cannot be directly applied to solve this problem. We propose two manifold proximal gradient descent ascent (MPGDA) algorithms for solving the problem. The first algorithm alternatively performs one or multiple manifold proximal gradient descent steps and a proximal ascent step at each iteration, and we prove that it can find an $\varepsilon$-game-stationary point and an $\varepsilon$-optimization-stationary point within $\mathcal{O}(\varepsilon^{-3})$ outer iterations. The second algorithm alternatively performs one manifold proximal gradient descent step and a proximal gradient ascent step, and we show that it can reach an $\varepsilon$-game-stationary point and an $\varepsilon$-optimization-stationary point within $\mathcal{O}(\varepsilon^{-4})$ outer iterations. Numerical experiments on an analytic example, fair sparse PCA, and sparse spectral clustering are conducted to illustrate the advantages of the proposed algorithms.
@@ 1 Introduction
where $\mathcal{M}$ is a Riemannian manifold embedded in a finite-dimensional Euclidean space $\mathbb{E}_{1}$, and $S$ is a convex set in another finite-dimensional Euclidean space $\mathbb{E}_{2}$, $f:\mathbb{E}_{1}\times\mathbb{E}_{2}\to\mathbb{R}$ is a continuously differentiable function and is concave with respect to y, $h:\mathbb{E}_{1}\to\mathbb{R}$ and $g:\mathbb{E}_{2}\to\mathbb{R}$ are some proper closed convex (possibly nonsmooth) functions. Throughout this paper, we adopt the following assumptions for problem (1).
@@
$(i)$ The manifold $\mathcal{M}$ and the convex set $S$ are compact. Let $\sigma_{y}:=\sup_{y\in S}\left\|y\right\|<\infty$.
@@
$(v)$ For all $u\in\mathbb{E}_{1},~w\in\mathbb{E}_{2}$ and $\alpha>0$, the following two proximal problems can be solved exactly and efficiently:
$\min_{x\in\mathbb{E}_{1}}h(x)+\frac{1}{2\alpha}\|x-u\|^{2},$ (2a)
$\min_{y\in S}g(y)+\frac{1}{2\alpha}\|y-w\|^{2}.$ (2b)
Problem (1) has many important applications in machine learning and signal processing. Note that if $f(x,y)$ is linear with respect to $y$ for fixed $x$, then $f(x,y)=\langle\mathcal{A}(x),y\rangle$ for some smooth operator $\mathcal{A}:\mathbb{E}_{1}\to\mathbb{E}_{2}$. Below, we present two representative examples and refer the interested readers to chenProximalGradientMethod2020; BriefIntroductiontoManifoldOptimization for more examples.
@@
Fair Sparse Principal Component Analysis (FSPCA). The classical PCA Hotelling1933Analysis is one of the most widely used dimensionality reduction techniques, which minimizes the total reconstruction error over the whole samples. However, in real-life applications, the data samples come from diverse classes and PCA may exhibit a higher reconstruction error for certain classes than the others. To address this disparity, the fair PCA is proposed samadiPriceFairPCA. Suppose the $m$ data samples belong to $n$ classes, and each class $i$ corresponds to $A_{i}\in\mathbb{R}^{m_{i}\times d}$ with $d$ being the dimension of the samples, $m_{i}$ being the number of samples in the $i$-th class, and $\sum_{i=1}^{n}m_{i}=m$. Then the fair PCA can be formulated as:
@@
where $r<d$ is the number of principal components, $\mathrm{Tr}(\cdot)$ represents the trace and $\mathrm{St}(d,r):=\{X\in\mathbb{R}^{d\times r}|X^{\top}X=I_{r}\}$ with $I_{r}$ denoting the $r\times r$ identity matrix. Recently, the fair sparse PCA is further proposed to promote sparsity in the principal components babuFairPrincipalComponent2023 as follows:
@@
where $\mu>0$ is a weighting parameter and $\left\|X\right\|_{1}:=\sum_{i,j}\left|x_{ij}\right|$ is the $\ell_{1}$-norm of the matrix $X$. In view of (Beck2017First, Example 4.10), problem (4) can be equivalently reformulated into
@@
Sparse Spectral Clustering (SSC). The SSC problem aims to partition $N$ data samples into $p$ groups such that similar data points are clustered together. In spectral clustering, a symmetric affinity matrix $W=[W_{ij}]_{N\times N}$ is constructed, where $W_{ij}\geq 0$ quantifies the pairwise similarity between samples $a_{i}$ and $a_{j}$. To promote sparsity and interpretability, the sparse spectral clustering framework is introduced in 2016ConvexSparseSpectralClustering, as follows:
@@
In recent years, nonconvex-concave (NC-C) minimax problems have received a tremendous amount of attention as they have found broad applications in diverse areas. Many studies focus on NC-C problems in the setting where both the constraints on $x$ and $y$ are closed convex sets and numerous methods have been developed, see, e.g., linGradientDescentAscent; linOptimalAlgorithmsMinimax; nouiehedSolvingClassNonConvex; ostrovskiiEfficientSearchFirstOrder2021; panEfficientAlgorithmNonconvexlinear2021a; 2022Zeroth; thekumparampilEfficientAlgorithmsSmooth; xuUnifiedSingleloopAlternating2023; yangFasterSingleloopAlgorithms; zhangSingleLoopSmoothedGradient. In contrast, there are relatively few works studying NC-C minimax problems in the setting where the constraint on $x$ is a Riemannian submanifold and the constraint on $y$ is a closed convex set. Among these works, Huang and Gao huangGradientDescentAscent2023 propose a Riemannian gradient descent-ascent algorithm for a smooth nonconvex-strongly concave (NC-SC) minimax problem and establish its iteration complexity of $\mathcal{O}(\varepsilon^{-2})$ for achieving $\varepsilon$-stationarity. Later, Xu et al. xuEfficientAlternatingRiemannian2024 present an alternating Riemannian/projected gradient descent-ascent (ARPGDA) algorithm with an iteration complexity of $\mathcal{O}(\varepsilon^{-3})$ for finding an $\varepsilon$-stationary point of smooth Riemannian nonconvex-linear (NC-L) minimax problem. Very recently, for problem (1) with additional assumptions that $h$ is smooth (but not necessarily convex) and $f(x,y)$ is linear with respect to $y$ for fixed $x$, the work xuRiemannianAlternatingDescent2024 extends the technique in heApproximationProximalGradient2024 and develops algorithms that can reach $\varepsilon$-stationary points within $\mathcal{O}(\varepsilon^{-3})$ iterations. It is worth noting that none of the methods in the aforementioned literature huangGradientDescentAscent2023; xuRiemannianAlternatingDescent2024; xuEfficientAlternatingRiemannian2024 can directly tackle problem (1) due to the following two reasons: (i) the convex function $h$ may be nonsmooth, and (ii) the function $f(x,y)$ is neither linear nor strongly concave with respect to $y$ for fixed $x$. Based on the needs of applications and the above background, it is highly desirable to develop efficient methods with theoretical convergence guarantees for solving the Riemannian nonsmooth NC-C minimax problem (1).
@@
The main contributions of this paper are summarized below. First, we introduce the notions of optimization stationarity as well as game stationarity for problem (1) and carefully study their relationships. Second, we propose two manifold proximal gradient descent ascent (MPGDA) algorithms for solving problem (1). The first algorithm, denoted by MPGDA-PA, performs one or multiple manifold proximal gradient descent steps to update $x$ and one proximal ascent step to update $y$ at each iteration. Note that the proximal ascent subproblem with respect to $y$ can generally be solved by FISTA (Beck2017First, Theorem 10.41) with a linear convergence rate. In particular, for the case of $f$ being linear with respect to $y$, addressing this subproblem reduces to solving a proximal problem in the form of (2b). The second algorithm, denoted by MPGDA-PGA, also performs one manifold proximal gradient descent step to update $x$ but uses a single proximal gradient ascent step to renew $y$ at each iteration. It is worth noting that the $y$-update in MPGDA-PGA requires only a single evaluation of the gradient $\nabla_{y}f$ and one step of solving the proximal problem (2b). This has much lower computational cost than that of MPGDA-PA in the case where $f$ is not linear with respect to $y$. Third, we prove that the MPGDA-PA algorithm can find an $\varepsilon$-game-stationary point of problem (1) within $\mathcal{O}(\varepsilon^{-3})$ outer iterations, while the MPGDA-PGA algorithm achieves this within $\mathcal{O}(\varepsilon^{-4})$ outer iterations. Finally, we illustrate the proposed algorithms via both analytic and simulated numerical examples.
@@
The rest of the paper is organized as follows. Section 2 presents preliminaries that will be used in this paper. Section 3 discusses the optimality condition of problem (1). Section 4 introduces the proposed MPGDA-PA method and establishes its iteration complexity. Section 5 introduces the proposed MPGDA-PGA method and establishes its iteration complexity. Section 6 reports the numerical experiments.
@@ 2 Preliminaries
In this section, we recall some basic notations and preliminary results which will be used in this paper. Let $\mathbb{E}$ represent a finite-dimensional Euclidean space with inner product $\langle\cdot,\cdot\rangle$ and the induced norm $\left\|\cdot\right\|$. For a linear operator $\mathcal{L}:\mathbb{E}_{1}\to\mathbb{E}_{2}$, let $\mathrm{rank}(\mathcal{L})$ denote the dimension of its image space $\{y\in\mathbb{E}_{2}\mid y=\mathcal{L}(x),~x\in\mathbb{E}_{1}\}$ and $\mathrm{ker}(\mathcal{L})$ denote its null space $\{x\in\mathbb{E}_{1}\mid\mathcal{L}(x)=0\}$ . Given a point $x\in\mathbb{E}$ and a set $C\subseteq\mathbb{E}$, let $\mathrm{dist}(x,C):=\inf_{y\in C}\|y-x\|$ and $\delta_{C}$ denote the indicator function associated with $C$. For a function $\varphi(x,y):\mathbb{E}_{1}\times\mathbb{E}_{2}\to\mathbb{R}$, let $\nabla_{x}\varphi(x,y)$ and $\nabla_{y}\varphi(x,y)$ respectively denote the partial gradients of $\varphi(x,y)$ with respect to $x$ and $y$ at the point $(x,y)$. Let $\mathcal{M}$ represent a Riemannian submanifold $\mathbb{E}_{1}$. For a function $\Phi:\mathcal{M}\to\mathbb{R}$, $\mathrm{grad}\Phi(x)$ denotes the Riemannian gradient of $\Phi$. For a function $\varphi:\mathcal{M}\times\mathbb{E}_{2}\to\mathbb{R}$, $\mathrm{grad}\varphi(\cdot,y)$ denotes the Riemannian gradient of $\varphi(\cdot,y)$ with $y$ fixed.
@@ 2.1 Riemannian submanifold
(boumal2023intromanifolds, Definition 3.10) Let $\mathbb{E}$ be a Euclidean space of dimension $d$. A nonempty subset $\mathcal{M}$ of $\mathbb{E}$ is called an embedded submanifold of $\mathbb{E}$ of dimension $n$ if either $n=d$ and $\mathcal{M}$ is open in $\mathbb{E}$ or n=dkn=d-k for some $k\geq 1$ and, for each $x\in\mathcal{M}$, there exists a neighborhood $U$ of $x$ in $\mathbb{E}$ and a smooth function $\psi:U\to\mathbb{R}^{k}$ such that:
@@
Let $\mathcal{M}$ be an embedded submanifold of $\mathbb{E}$. The tangent space $T_{x}\mathcal{M}$ to $\mathcal{M}$ at $x\in\mathcal{M}$ is identified with $\mathbb{E}$ when $\mathcal{M}$ is open in $\mathbb{E}$; otherwise, $T_{x}\mathcal{M}=\mathrm{ker}(\mathrm{D}\psi(x))$ , where $\psi$ is a local defining function for $\mathcal{M}$ at $x$ (boumal2023intromanifolds, Theorem 3.15). The embedded submanifold $\mathcal{M}$ is called a Riemannian submanifold of $\mathbb{E}$, when equipped with the Riemannian metric induced by the ambient space $\mathbb{E}$. Namely, for each $x\in\mathcal{M}$, the inner product on $T_{x}\mathcal{M}$ is defined as $\langle u,v\rangle_{x}:=\langle u,v\rangle$ for $u,~v\in T_{x}\mathcal{M}$. The Riemannian gradient of a smooth function $\varphi:\mathbb{E}\to\mathbb{R}$ at a point $x\in\mathcal{M}$ is given by $\mathrm{grad}\varphi(x):=\mathrm{Proj}_{T_{x}\mathcal{M}}(\nabla\varphi(x))$, where $\mathrm{Proj}_{T_{x}\mathcal{M}}(\cdot)$ is the Euclidean projection operator onto $T_{x}\mathcal{M}$. The normal space $N_{x}\mathcal{M}$ of a Riemannian submanifold $\mathcal{M}$ of $\mathbb{E}$ is the orthogonal complement of the tangent space $T_{x}\mathcal{M}$ in $\mathbb{E}$ (boumal2023intromanifolds, Definition 5.48).
The tangent bundle of $\mathcal{M}$ is the disjoint union of its tangent spaces : $T\mathcal{M}:=\{(x,v)\mid x\in\mathcal{M},~v\in T_{x}\mathcal{M}\}$. A retraction on a manifold $\mathcal{M}$ is a smooth mapping $R$ from the tangent bundle $T\mathcal{M}$ onto $\mathcal{M}$ such that, for any $(x,v)\in T_{x}\mathcal{M}$, its restriction $R_{x}:T_{x}\mathcal{M}\to\mathcal{M}$ satisfies (i) $R_{x}(0)=x$ and (ii) $\mathrm{D}R_{x}(0)[v]=v$ (boumal2023intromanifolds, Definition 3.47).
The concept of a retraction offers a practical link between manifold and its tangent bundle. The following properties of retractions are needed for convergence analysis in this paper.
@@
(boumalGlobalRatesConvergence2019, Appendix B) Let $\mathcal{M}$ be a compact Riemannian submanifold of Euclidean space $\mathbb{E}$ equipped with norm $\left\|\cdot\right\|$. $R:T\mathcal{M}\to\mathcal{M}$ is a retraction. Then there exist constants $M_{1}>0$ and $M_{2}>0$ such that the following two inequalities hold for any $x\in\mathcal{M}$ and $v\in T_{x}{\mathcal{M}}$:
@@ 2.2 Generalized subdifferential
For a function $\varphi:\mathbb{E}\to\mathbb{R}$, the usual one-sided directional derivative of $\varphi$ at $x\in\mathbb{E}$ along the direction $v\in\mathbb{E}$ is defined by
@@
In contrast to the usual one-sided directional derivative, the (Clarke) generalized directional derivative does not presuppose the existence of any limit (1983Optimization, Section 2.1). Specifically, the (Clarke) generalized directional derivative of $\varphi$ at $x\in\mathbb{E}$ along the direction $v\in\mathbb{E}$ is defined as
@@
If $\varphi:\mathbb{E}\to\mathbb{R}$ is convex, then for any $x\in\mathbb{E}$, $\partial f(x)$ coincides with the subdifferential at $x$ in the sense of convex analysis.
A function $\varphi:\mathbb{E}\to\mathbb{R}$ is said to be regular at $x\in\mathbb{E}$, provided that for all $v\in\mathbb{E}$, the usual one-sided directional derivative $\varphi^{\prime}(x;v)$ exists and $\varphi^{\prime}(x;v)=\varphi^{\circ}(x;v)$ (1983Optimization, Definition 2.3.4). If $\varphi_{1}:\mathbb{E}\to\mathbb{R}$ and $\varphi_{2}:\mathbb{E}\to\mathbb{R}$ are locally Lipschitz continuous and regular at $x\in\mathbb{E}$, then $\varphi_{1}+\varphi_{2}$ are regular at $x$ and $\partial(\varphi_{1}+\varphi_{2})(x)=\partial\varphi_{1}(x)+\partial\varphi_{2}(x)$ (1983Optimization, Proposition 2.3.3, Corollary 3, Proposition 2.3.6). If $\varphi:\mathbb{E}\to\mathbb{R}$ is a continuously differentiable function or a real-valued convex function, then it is regualr at any $x\in\mathbb{E}$ (1983Optimization, Proposition 2.3.6). A function $\varphi:\mathbb{E}\to\mathbb{R}$ is said to be $\ell$-weakly convex if $\varphi(\cdot)+(\ell/2)\|\cdot\|^{2}$ is convex. It follows that the real-valued weakly convex function $\varphi$ is regular at any $x\in\mathbb{E}$. In addition, the indication function $\delta_{\mathcal{M}}(x)$ of a Remannian submanifold $\mathcal{M}$ in $\mathbb{E}$ is regular at any $x\in\mathcal{M}$ (1983Optimization, Theorem 2.4.7) and $\partial\delta_{\mathcal{M}}(x)=N_{x}\mathcal{M}$ (1983Optimization, Proposition 2.4.12).